# Mixed Signals: Understanding Model Disagreement in Multimodal Empathy Detection

### v1

@@ Appendix A Model Training Details
A.1 Unimodal Model Training Details
Each model is trained on a binary empathy classification task using precomputed 768-dimensional embeddings. We freeze all but the final two transformer layers and train for 15 epochs with a learning rate of 5e-6 and batch size of 8.
A.2 Fusion Model Details
Each unimodal model representation is independently gated and passed through an attention mechanism that computes modality-specific weights. The weighted embeddings are aggregated and classified using a three-layer feedforward network with max pooling. The fusion model is trained for 10 epochs using a learning rate of 1e-4 and includes modality dropout during training.
Appendix B Dataset Details
We use a multimodal empathy dataset Chen et al. (2024) consisting of 346 English-language videos totaling approximately 53 hours, collected from YouTube between 2020 and 2022 using keywords like “empathy” and “empathetic training.” The dataset includes empathy training sessions, therapy roleplays, interviews, TED Talks, and TV/movie scenes, comprising both acted (62%) and spontaneous (38%) speech. Each video was labeled by at least three expert annotators as either empathetic or neutral, with final labels determined by majority vote. Metadata such as speaker gender, topic, and emotional context was manually annotated, covering themes like therapy, parenting, workplace dynamics, and social relationships. From this collection, a subset of 65 videos was transcribed, diarized, and manually re-aligned using Praat to ensure accurate speaker segmentation and time alignment. This process resulted in 1,718 annotated segments with speaker labels, timestamps, transcripts, and empathy stage annotations, enabling fine-grained analysis of empathy in naturalistic and semi-scripted settings.
@@ Appendix C Annotation Instructions
Annotators were asked to provide two judgments per example, labeling each as either empathetic or neutral (Figure 4). A excerpt describing empathy (drawn from the Encyclopedia of Social Psychology, Volume 1, Baumeister and Vohs (2007)) was provided to ensure a consistent conceptual foundation for annotation:
@@ Appendix D Full Feature Comparisons
Tables 6, 7 and 8 provide additional results from the t-tests comparing examples across different confidence quadrants. Table 6 provides an internal comparison between the disagreement quadrants. Table 7 presents the full version of the audio feature comparisons summarized in Table 3. Table 8 expands on the facial feature comparisons shown in Table 4.
@@
AU04: Brow Lowerer 0.0106 red $>$ blue 0.3682 green $>$ blue
AU12: Lip Corner Puller 0.0174 blue $>$ red 0.8977 green $>$ blue
AU05: Upper Lid Raiser 0.1837 blue $>$ red <0.0001 blue $>$ green
AU17: Chin Raiser 0.2256 red $>$ blue 0.9802 blue $>$ green
AU10: Upper Lip Raiser 0.2275 blue $>$ red 0.6700 green $>$ blue
AU45: Blink 0.3200 blue $>$ red 0.7462 green $>$ blue
AU07: Lid Tightener 0.3252 blue $>$ red 0.9318 blue $>$ green
AU14: Dimpler 0.4593 red $>$ blue 0.0652 green $>$ blue
AU20: Lip Stretcher 0.5701 blue $>$ red 0.7907 blue $>$ green
AU09: Nose Wrinkler 0.6211 blue $>$ red 0.7639 green $>$ blue
AU25: Lips Part 0.6227 blue $>$ red 0.7492 blue $>$ green
AU01: Inner Brow Raiser 0.6529 blue $>$ red 0.4674 green $>$ blue
AU23: Lip Tightener 0.6630 red $>$ blue 0.3474 green $>$ blue
AU28: Lip Suck 0.6735 red $>$ blue 0.9846 green $>$ blue
AU26: Jaw Drop 0.6851 red $>$ blue 0.4596 blue $>$ green
AU06: Cheek Raiser 0.7097 blue $>$ red 0.3201 green $>$ blue
AU15: Lip Corner Depressor 0.9528 red $>$ blue 0.4834 green $>$ blue
AU02: Outer Brow Raiser 0.9647 blue $>$ red 0.6677 green $>$ blue
@@ Appendix E Feature Distributions
Figures 5 and 6 visualize the distributions of key features across confidence quadrants. Figure 5 presents the distribution of selected audio features (e.g., pitch, intensity) for red, green, and blue examples, highlighting acoustic patterns associated with model disagreement. Figure 6 shows activation rates for facial Action Units (AUs) in red, green, and blue examples, illustrating how specific facial expressions vary across agreement conditions. These visualizations complement the statistical comparisons reported in Tables 7 and 8, providing a more interpretable view of the underlying feature dynamics.
@@
Figure 6: AU activation rates for red, green, and blue examples. Red bars indicate examples where the unimodal visual model predicted correctly but the multimodal model did not (Red: Unimodal $>$ 0.5, Multimodal $<$ 0.5). Green bars show the reverse. Blue bars indicate examples where both the unimodal and multimodal models correctly predicted the label.

### v3

@@ Appendix B Model Training Details
B.1 Unimodal Model Training Details
Each model is trained on a binary empathy classification task using precomputed 768-dimensional embeddings. We freeze all but the final two transformer layers and train for fifteen epochs with a learning rate of 5e-6 and batch size of eight.
B.2 Fusion Model Details
Each unimodal model representation is independently gated and passed through an additive attention mechanism that computes modality-specific weights. The weighted embeddings are aggregated and classified using a three-layer feedforward network with max pooling. The fusion model is trained for ten epochs using a learning rate of 1e-4 and includes modality dropout during training. To characterize how the model balances each of the three modalities at inference time, we computed the per-modality gate-weight distributions over the full test set (Table  9). The mean gate weights indicate that our model allocates substantial importance to each modality, with only a slight preference toward audio and video. The high variance also shows that the model dynamically adapts its reliance on each modality on a per‐sample basis. Thus, our fusion model draws substantially on all three streams; no single modality is systematically favored.
Modality Mean Standard Deviation
Text 0.430 0.297
Audio 0.502 0.227
Video 0.476 0.315
Table 9: Per-modality gate-weight distributions over the full test set.
@@ Appendix C Annotation Instructions
Annotators were asked to provide two judgments per example, labeling each as either empathetic or neutral (Figure 4). A excerpt describing empathy (drawn from the Encyclopedia of Social Psychology, Volume 1, baumeister) was provided to ensure a consistent conceptual foundation for annotation:
@@ Appendix D Full Feature Comparisons
Tables 10, 11 and 12 provide additional results from the t-tests comparing examples across different confidence quadrants. Table 10 provides an internal comparison between the disagreement quadrants. Table 11 presents the full version of the audio feature comparisons summarized in Table 3. Table 12 expands on the facial feature comparisons shown in Table 4.
@@
AU04: Brow Lowerer 0.0106 red >> blue 0.3682 green >> blue
AU12: Lip Corner Puller 0.0174 blue >> red 0.8977 green >> blue
AU05: Upper Lid Raiser 0.1837 blue >> red <0.0001 blue >> green
AU17: Chin Raiser 0.2256 red >> blue 0.9802 blue >> green
AU10: Upper Lip Raiser 0.2275 blue >> red 0.6700 green >> blue
AU45: Blink 0.3200 blue >> red 0.7462 green >> blue
AU07: Lid Tightener 0.3252 blue >> red 0.9318 blue >> green
AU14: Dimpler 0.4593 red >> blue 0.0652 green >> blue
AU20: Lip Stretcher 0.5701 blue >> red 0.7907 blue >> green
AU09: Nose Wrinkler 0.6211 blue >> red 0.7639 green >> blue
AU25: Lips Part 0.6227 blue >> red 0.7492 blue >> green
AU01: Inner Brow Raiser 0.6529 blue >> red 0.4674 green >> blue
AU23: Lip Tightener 0.6630 red >> blue 0.3474 green >> blue
AU28: Lip Suck 0.6735 red >> blue 0.9846 green >> blue
AU26: Jaw Drop 0.6851 red >> blue 0.4596 blue >> green
AU06: Cheek Raiser 0.7097 blue >> red 0.3201 green >> blue
AU15: Lip Corner Depressor 0.9528 red >> blue 0.4834 green >> blue
AU02: Outer Brow Raiser 0.9647 blue >> red 0.6677 green >> blue
@@ Appendix E Feature Distributions
Figures 5 and 6 visualize the distributions of key features across confidence quadrants. Figure 5 presents the distribution of selected audio features (e.g., pitch, intensity) for red, green, and blue examples, highlighting acoustic patterns associated with model disagreement. Figure 6 shows activation rates for facial Action Units (AUs) in red, green, and blue examples, illustrating how specific facial expressions vary across agreement conditions. These visualizations complement the statistical comparisons reported in Tables 11 and 12, providing a more interpretable view of the underlying feature dynamics.
@@
Figure 6: AU activation rates for red, green, and blue examples. Red bars indicate examples where the unimodal visual model predicted correctly but the multimodal model did not (Red: Unimodal >> 0.5, Multimodal << 0.5). Green bars show the reverse. Blue bars indicate examples where both the unimodal and multimodal models correctly predicted the label.
Appendix F Quadrant Examples
Table 13 shows video frames, transcripts, annotator judgments, and the true labels for examples from each confidence plot quadrant. These examples illustrate that disagreement quadrants often contain more ambiguous instances for both humans and models where cues from different modalities may conflict, while examples from agreement quadrants typically display alignment between modalities.
[Uncaptioned image] Transcript: "I assume you don’t know who emailed me for the emergency sessions" Quadrant: Red True Label: Empathetic Annotator 1: Neutral Annotator 2: Empathetic [Uncaptioned image] Transcript: "In fact research suggests we spend about 55 percent of our day…" Quadrant: Blue True Label: Neutral Annotator 1: Neutral Annotator 2: Neutral
[Uncaptioned image] Transcript: "One of the reasons I wanted to come here tonight was to discuss our future." Quadrant: Yellow True Label: Neutral Annotator 1: Empathetic Annotator 2: Empathetic [Uncaptioned image] Transcript: "It’s good to have you here um especially to talk about a topic that i think is one of the more sensitive topics that we we’re discussing in society today…" Quadrant: Green True Label: Empathetic Annotator 1: Neutral Annotator 2: Empathetic
Table 13: Example clips from each disagreement quadrant with transcript and labels.